OpenStudy (anonymous):

Which expression is equivalent to the area of metal sheet required to make this square-shaped traffic sign? A square shaped traffic sign is shown with length of one side labeled as x + 1. x2 + 2x + 1 x2 + x + 1 x2 + 2x x2 + 1

OpenStudy (yttrium):

Do you know the formula for solving for the area of a square?

OpenStudy (anonymous):

nope

OpenStudy (yttrium):

The formula for that is A = s^2, where A is the area and s is the length of the side.

OpenStudy (anonymous):

still dont get it

OpenStudy (yttrium):

In order to solve for the are of square, you just have to square the length of its side. And that's it. For example, you have a square with side=4, then the area is 16square units.

OpenStudy (anonymous):

still dont get it

OpenStudy (yttrium):

side = 4 solving for area, you have to square it. meaning, you have to times the side to itself. Therefore, A = 4*4 = 16

OpenStudy (yttrium):

In your case, the side is equal to x+1. Therefore, what are you going to do?

OpenStudy (anonymous):

\[x ^{2}+1\]

OpenStudy (anonymous):

is that the answer

OpenStudy (ybarrap):